A company has one year, zero-coupon debt outstanding with face value $83 million. The company's current market value of equity is $100 million. The firm will experience a free cash flow at year 1 and no further free cash flows. The free cash flow will be $225 million with probability 70% and $130 million with probability 30%. What is the equity cost of capital?
Answer #1

Expected free cash flow = 0.7*225 + 0.3*130

=196.50 million

From this 83 million has to be given back due to Zero coupon bond expiry

Total cash flow available for equity=196.50 - 83 = $113.50

Equity cost of capital is the required return that the firm will create in 1 year = (113.50-100)/100 =13.50%